Compute the density (in units of m3

) of an ideal gas under the following

conditions:

(a) At 0 C and 760 Torr pressure (1 Torr ¼ 1 mmHg). This is called the

Loschmidt number.

(b) In a vacuum of 103

Torr at room temperature (20 C). This number is a

useful one for the experimentalist to know by heart (103

Torr ¼ 1 μ).


1
Expert's answer
2021-06-14T12:19:41-0400

According to ideal gas law:


"n=\frac{p}{k_BT}."

(a) In this scenario, at 760 torr and 0°C:


"n=\frac{760·133.322}{(1.38·10^{-23})273.15}=2.687·10^{25}\text{ m}^{-3}."

(b) At room temperature:


"n=\frac{(10^{-3})·133.322}{(1.38·10^{-23})(273.15+20)}=3.294·10^{19}\text{ m}^{-3}."